# Autonomous Stabilization of Floquet States Using Static Dissipation

### v1

@@ I Introduction
The ability to experimentally control and manipulate quantum systems has rapidly advanced over the last two decades. One indispensable tool in the quantum mechanic’s toolkit is Floquet engineering, in which the properties of the system are qualitatively modified by strong periodic driving. It underlies the creation of optical lattices and synthetic band structures [1, 2, 3, 4, 5, 6], artificial gauge fields [7, 8, 2, 9, 10], topological phases impossible in static systems [11, 12, 13, 14, 15], topological charge and photon pumps [16, 17, 18, 19, 20, 21, 22, 23, 24], and exotic quantum codes [25], to name a few.
However, Floquet engineering is always limited by heating. There are three conceptually distinct mechanisms for heating. Consider a single degree of freedom with a few energy levels subject to Floquet driving. The stationary states of the system (known as quasienergy states) involve admixtures of both ground and excited states. Accounting for dissipation and dephasing in the laboratory frame leads to both excitation and de-excitation processes in the quasienergy state basis, and thus hot steady states in this basis. Second, restricting to a few levels is always an approximation in a physical system. Higher order processes often allow the system to absorb energy from the drive and leak out of the intended operating subspace. Third, with many interacting degrees of freedom, the system can absorb energy from the drive and heat up, even if the first two mechanisms are suppressed [26, 27, 28].
@@
Conventional dissipative preparation methods leverage native decay processes to successfully cool into ground states [29, 30] or dark states [31]. However, in the presence of strong drives, laboratory frame decay processes cause both excitation and de-excitation. This, combined with the time-dependence of the states, causes conventional autonomous cooling schemes like optical pumping [29] and dark-state engineering [32, 31, 33] to fail. Their generalization to the time-dependent case is not known.
One route to stabilizing more general, but still time-independent, states is to engineer new dissipation terms distinct from the native dissipators [34, 33]. These methods have been used to stabilize arbitrary static single-qubit states [35, 36, 37], bosonic Kerr-Cat qubits [38, 39, 40], code spaces in a superconducting-qubit Cat code [41], and entangled states between multiple qubits [42, 43, 44, 45, 46, 47, 48, 49]. In principle, periodic modulation of such experimentally controlled dissipators could implement spontaneous emission into a specific quasienergy state. Such an approach is however highly fine-tuned, requiring knowledge of the target quasienergy state at all times during the drive cycle.
A different route to stabilization is to actively extract heat from the system continuously [50, 51, 52, 53] or stroboscopically [54, 55, 56] into an auxiliary system. This type of approach has been extended to Floquet systems in the limit of high-frequency driving. Ref. [53] proposed stabilization of the computational subspace of a Kapitzonium qubit using auxiliary oscillators, while Ref. [56] cooled into the ground state of a static many-qubit Hamiltonian by cycling between a Trotterized Hamiltonian evolution step and an auxiliary qubit reset step.
@@
As an application beyond two-level systems, we develop the theory of a topological photon pump [18, 19, 20, 21, 22] coupled to an auxiliary lossy cavity. The topological photon pump is a strongly driven cavity-QED system with a long-lived photon current of topological origin. We show that the auxiliary cavity brings the value of the photon current closer to its ideal value, and increases the fidelity of high-photon number Fock states that are prepared in the cavity at special times [22].
@@
The remainder of the manuscript is organized as follows. In Sec. II, we describe the physical setup involving a periodically modulated spin coupled to a lossy cavity, which showcases the autonomous stabilization effect. In Sec. III, we present experimental evidence for stabilization in this physical setup in the low-frequency (adiabatic) regime. We turn to the general theory in Sec. IV, and provide numerical evidence for stabilization in all frequency regimes. Lastly, in Sec. V, we show that dissipative stabilization enhances the performance of the topological photon pump [20, 22].
@@ II The Physical Set-up
with a fixed modulation frequency $\omega_{\mathrm{mod}}$ and constant magnetic field amplitude $B_{0}$. The spin is subject to additional external dissipation, which we return to below. See the schematic in Fig. 1(a).
@@ III Experimental Observation of Stabilization in the Adiabatic Regime
In this section, we present an experimental implementation of the model in Eq. (3) in the adiabatic limit and demonstrate stabilization of a quasienergy state using a transmon qubit [57, 58] coupled to a superconducting cavity. We show that, when the resonance condition in Fig. 1(b) holds, the transmon is stabilized to a chosen instantaneous eigenstate for timescales that far exceed the native $T_{1}$ and $T_{2}$ decoherence times of the qubit.
@@
To create an effective rotating magnetic field of the form in $H_{B}$ [Eq. (1)], we combine simultaneous frequency modulation of the $g-e$ transition of the transmon and strong microwave driving at the mean $g-e$ transition frequency [22]. The flux bias sets the frequency of the $g-e$ transition [57, 58], and therefore creates an effective $B_{z}$. On the other hand, the Rabi rate induced by a constant AC microwave drive $A\cos(\omega_{d}t)$ implements an effective static $B_{x}$ term in a frame rotating at the drive frequency $\omega_{d}$ [59]. To synthesize $H_{B}$, we first select an operating point using a DC flux bias $\Phi_{\mathrm{ext}}=\Phi_{\mathrm{static}}$, then set the microwave drive frequency to be resonant with the qubit transition at this bias point, and operate in a rotating frame at this frequency. Rotation of the effective field at frequency $\omega_{\mathrm{mod}}$ is induced by modulating both the flux bias $[\Phi_{\mathrm{ext}}=\Phi_{\mathrm{static}}+\delta\Phi\sin(\omega_{\mathrm{mod}}t)]$ and the AC drive $[V_{\mathrm{drive}}=A\cos(\omega_{\mathrm{mod}}t)\cos(\omega_{d}t)]$ in quadrature. See Appendix B for more details on the effective rotating field. In order to be deep within the adiabatic regime, we use $B_{0}=2\pi\times 80$ MHz, and modulation frequencies of $\omega_{\mathrm{mod}}=2\pi\times 0.75$ MHz.
@@ V Dissipative Stabilization of a Topological Photon Pump
Quantum states can serve as resources for classically difficult tasks, making reliable and stable preparation of these states a useful module in a variety of quantum engineering protocols. The most familiar examples of such resource states are static—a given scheme would call for the preparation of a state |$\ket{\psi_{\mathrm{resource}}}$, which does not evolve except when acted on by the protocol [60, 61, 62, 63, 64]. However, some schemes explicitly call for time-varying periodic states, |$\ket{\psi_{\mathrm{resource}}(t)}$, and utilize this time dependence as part of their dynamical protocol [18, 19, 20, 21, 22]. These periodic resource states can be stabilized by our methods.
Cavity state boosting [22] is a topological pumping protocol which can be used to coherently translate, or boost, the quantum state of a cavity in the Fock basis. This enables the preparation of highly excited non-classical states, such as Fock states, if lower photon number states can be reliably prepared (see Refs. [65, 66, 67, 68, 69] for other methods to prepare Fock states). More generally, given an initial cavity state written in terms of Fock states
@@
The required resource state is a strongly varying qubit state that covers the Bloch sphere [18, 19]. When this state is chosen correctly and coupled to the cavity with exchange (Jaynes-Cummings) interactions, the influence of the qubit on the cavity is to effectively implement a Thouless pump in photon number space. On average, every cycle of the qubit state pumps a integer number $C$ of photons into the cavity, leading to an average increase in the photon number with the quantized rate
@@
where $\dot{n}=\mathrm{d}n/\mathrm{d}t$ is the average photon number current into the cavity [20]. Henceforth, we take $C=1$. Further, Ref. [22] showed that at times $t^{*}$ which are close to being a common period of the cavity and the qubit, $t^{*}\approx N(2\pi/\Delta_{b})\approx MT_{\mathrm{mod}}$, the cavity state rephases into a boosted copy of the initial state [Eq. (12)]. The cavity frequency, $\Delta_{b}$, is defined in Eq. (14).
Reference [22] appealed to an adiabatic control scheme for the resource qubit. By implementing a strong, slow drive on the qubit, the back-action of the cavity on the qubit is negligible except when they need to exchange energy. This necessitates an operating regime where the pumping rate—given by $1/T_{\mathrm{mod}}$—is very low, which in turn limits the size of a cavity state that can be prepared in a time- or loss-limited experiment. The introduction of an auxiliary lossy element can stabilize the required resource state, lending additional robustness to the boosting protocol and hence allowing for faster, more reliable preparation of excited cavity states.
@@
Cavity state boosting (and topological photon pumping more broadly [20]) requires the boost cavity energy and drive frequencies to be much smaller than the qubit energy and interaction: $\Delta_{\mathrm{b}},\omega_{\mathrm{mod}}\ll B_{0},g_{\mathrm{b}}\sqrt{n_{\mathrm{b}}}$. Meanwhile, stabilization requires a resonance condition between the qubit quasienergy (which can be replaced by the instantaneous energy near the adiabatic limit) and the stabilizing cavity: $|\Delta_{\mathrm{s}}-B_{0}|\lesssim g_{\mathrm{s}}$. In order for the boost cavity to not couple strongly to the stabilizing cavity, and inherit its photon loss, we also demand that they are far detuned: $|\Delta_{\mathrm{s}}-\Delta_{\mathrm{b}}|\gg g_{\mathrm{b}},g_{\mathrm{s}}$.
@@ VI Conclusion and Outlook
However, there are two aspects of our results that suggest that the same scheme can combat these other kinds of heating. First, our experimental implementation is carried out on a transmon qubit, which is not a true two-level system. In fact, the transmon is notorious for being only loosely describable as a two-level system due to its relatively low anharmonicity of $100-250$ MHz (or roughly $2\%-5\%$). The data in Fig. 2 shows persistent oscillations for $100\ \mu$s and over $100$ drive periods without signs of uncompensated heating due to the higher levels of the transmon, despite applied drive strengths which are roughly one third of the anharmonicity. Second, the stabilized topological photon pump described in Sec. V uses a second lossy cavity to stabilize a Floquet system consisting of both a qubit and a cavity. The presence of autonomous stabilization in both of these cases despite the larger Hilbert space is a strong indicator that it is possible to suppress Floquet heating in general using our simple scheme, rather than the contributions of excited-state lifetime alone. We also have evidence that the scheme can combat heating due to many-body interactions, and can cool into desired ground states of effective Floquet Hamiltonians [70]. However, a full characterization of the capability of autonomous stabilization with a static auxiliary remains an outstanding question.
Optimal stabilization, with the fastest possible stabilization time and highest possible steady-state purity, relies on exact resonances in the augmented spin-cavity system. However, significant stabilization can be achieved for a much wider range of parameters. Supplementary experimental data, shown in Appendix F, shows that significant stabilization is observed for all configurations in the range $|\delta|\lesssim g$. As the transmon decay rate is only a fourth of the photon decay rate, using an even lossier cavity (up to $\kappa\approx g$) would result in stabilization for an even wider range of parameters. Additionally, we find that elliptical drive fields, which are only resonant during part of the drive cycle, also exhibit stabilization. The combination of these two experimental observations is a strong indicator that a single lossy auxiliary degree of freedom can be used to counteract heating into all excited states within a range of roughly $g$ of the resonance, and hence systematically remove population from entire targeted Floquet bands. Alternatively, an auxiliary consisting of a band-pass filter, rather than a single cavity, could be used to enhance loss for a large number of Floquet states, as was recently proposed in Ref. [53] for a Kaptizonium qubit in the high-frequency Floquet limit. Implementations along these lines would give the ability to, for example, cool directly into the ground band of a Floquet engineered lattice, rather than relying on adiabatic band mapping, which is generally only possible in synthetic lattice systems.

### v2

@@ I Introduction
The ability to experimentally control and manipulate quantum systems has rapidly advanced over the last two decades. One indispensable tool in the quantum mechanic’s toolkit is Floquet engineering, in which the properties of the system are qualitatively modified by strong periodic driving. It underlies the creation of optical lattices and synthetic band structures Bloch et al. (2008); Goldman and Dalibard (2014); Bukov et al. (2015); Oka and Aoki (2009); Rechtsman et al. (2013); Holthaus (2015), artificial gauge fields Spielman (2009); Lin et al. (2009); Goldman and Dalibard (2014); Roushan et al. (2017); Rosen et al. (2024), topological phases impossible in static systems Rudner et al. (2013); Harper et al. (2020); Wintersperger et al. (2020); Khemani et al. (2019); Else et al. (2020), topological charge and photon pumps Thouless (1983); Citro and Aidelsburger (2023); Martin et al. (2017); Crowley et al. (2019); Nathan et al. (2019); Psaroudaki and Refael (2021); Long et al. (2022, 2021); Nathan et al. (2021), and exotic quantum codes Hastings and Haah (2021), to name a few.
However, Floquet engineering is always limited by heating. There are three conceptually distinct mechanisms for heating. Consider a single degree of freedom with a few energy levels subject to Floquet driving. The stationary states of the system (known as quasienergy states) involve admixtures of both ground and excited states. Accounting for dissipation and dephasing in the laboratory frame leads to both excitation and de-excitation processes in the quasienergy state basis, and thus hot steady states in this basis. Second, restricting to a few levels is always an approximation in a physical system. Higher order processes often allow the system to absorb energy from the drive and leak out of the intended operating subspace. Third, with many interacting degrees of freedom, the system can absorb energy from the drive and heat up, even if the first two mechanisms are suppressed Lazarides et al. (2014); Ponte et al. (2015); Mori et al. (2018).
@@
Conventional dissipative preparation methods leverage native decay processes to successfully cool into ground states Happer (1972); Phillips (1998) or dark states Fleischhauer et al. (2005). However, in the presence of strong drives, laboratory frame decay processes cause both excitation and de-excitation. This, combined with the unboundedness of Floquet spectra from above and below, causes conventional autonomous cooling schemes like optical pumping Happer (1972) and dark-state engineering Diehl et al. (2008); Fleischhauer et al. (2005); Harrington et al. (2022) to fail. Their generalization to the time-dependent case is not known.
One route to stabilizing more general, but still time-independent, states is to engineer new dissipation terms distinct from the native dissipators Poyatos et al. (1996); Harrington et al. (2022). For example, in superconducting qubit architectures, two photon loss has been used to stabilize bosonic Kerr-Cat qubits Leghtas et al. (2015); Puri et al. (2017); Grimm et al. (2020), while selective photon addition can stabilize code spaces in a Cat code Gertler et al. (2021). Collective dissipators can also stabilize target entangled states of atoms Plenio et al. (1999); Kraus et al. (2008); Verstraete et al. (2009); Barreiro et al. (2011); Lin et al. (2013); Shankar et al. (2013); Brown et al. (2022). In principle, periodic modulation of engineered dissipators could implement spontaneous emission into a specific quasienergy state. Such an approach is however highly fine-tuned, requiring knowledge of the target quasienergy state at all times during the drive cycle.
A different route to stabilization is to actively extract heat from the system continuously Home et al. (2009a, b); Raghunandan et al. (2020); Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023); Murch et al. (2012); Lu et al. (2017); Ma et al. (2019); Li et al. (2024) or stroboscopically Lloyd (1996); Sarovar and Milburn (2005); Mi et al. (2024) into an auxiliary system. This type of approach has been used to stabilize arbitrary static single-qubit states Murch et al. (2012); Lu et al. (2017); Ma et al. (2019) and entangled states of two qubits Li et al. (2024). These methods have also been extended to Floquet systems in the limit of high-frequency driving, where they stabilize the ground state space of static effective Hamiltonians. Refs. Petiziol and Eckardt (2022); Wang and Safavi-Naeini (2023) numerically demonstrated such stabilization, while Ref. Mi et al. (2024) experimentally cooled into the ground state of an effective many-qubit Hamiltonian by cycling between a Trotterized Hamiltonian evolution step and an auxiliary qubit reset step.
@@
As an application beyond two-level systems, we develop the theory of a topological photon pump Martin et al. (2017); Crowley et al. (2019); Nathan et al. (2019); Psaroudaki and Refael (2021); Long et al. (2022) coupled to an auxiliary lossy cavity. The topological photon pump is a strongly driven cavity-QED system with a long-lived photon current of topological origin. We show that the auxiliary cavity brings the value of the photon current closer to its ideal value, and increases the fidelity of high-photon number Fock states that are prepared in the cavity at special times Long et al. (2022).
@@
The remainder of the manuscript is organized as follows. In Sec. II, we describe the physical setup involving a periodically modulated spin coupled to a lossy cavity, which showcases the autonomous stabilization effect. In Sec. III, we present experimental evidence for stabilization in this physical setup in the low-frequency (adiabatic) regime. We turn to the general theory in Sec. IV, and provide numerical evidence for stabilization in all frequency regimes. Lastly, in Sec. V, we show that dissipative stabilization enhances the performance of the topological photon pump Nathan et al. (2019); Long et al. (2022).
@@ II The Physical Set-up
with a fixed modulation frequency $\omega_{\mathrm{mod}}$ and constant magnetic field amplitude $B_{0}$. The model happens to be soluble (it reduces to a static problem in the frame rotating about the $y$-axis with frequency $\omega_{\mathrm{mod}}$). This solubility is however irrelevant to what follows, as the augmented system with dissipation is not soluble. See the schematic in Fig. 1(a).
@@ III Experimental Observation of Stabilization in the Adiabatic Regime
In this section, we present an experimental implementation of the model in Eq. (3) in the adiabatic limit and demonstrate stabilization of a quasienergy state using a transmon qubit Blais et al. (2021); Krantz et al. (2019) coupled to a superconducting cavity. We show that, when the resonance condition in Fig. 1(b) holds, the transmon is stabilized to a chosen instantaneous eigenstate for timescales that far exceed the native $T_{1}$ and $T_{2}$ decoherence times of the qubit.
@@
To create an effective rotating magnetic field of the form in $H_{B}$ [Eq. (1)], we combine simultaneous frequency modulation of the $g-e$ transition of the transmon and strong microwave driving at the mean $g-e$ transition frequency Long et al. (2022). The flux bias sets the frequency of the $g-e$ transition Blais et al. (2021); Krantz et al. (2019), and therefore creates an effective $B_{z}$. On the other hand, the Rabi rate induced by a constant AC microwave drive $A\cos(\omega_{d}t)$ implements an effective static $B_{x}$ term in a frame rotating at the drive frequency $\omega_{d}$ Walls and Milburn (2008). To synthesize $H_{B}$, we first select an operating point using a DC flux bias $\Phi_{\mathrm{ext}}=\Phi_{\mathrm{static}}$, then set the microwave drive frequency to be resonant with the qubit transition at this bias point, and operate in a rotating frame at this frequency. Rotation of the effective field at frequency $\omega_{\mathrm{mod}}$ is induced by modulating both the flux bias $[\Phi_{\mathrm{ext}}=\Phi_{\mathrm{static}}+\delta\Phi\sin(\omega_{\mathrm{mod}}t)]$ and the AC drive $[V_{\mathrm{drive}}=A\cos(\omega_{\mathrm{mod}}t)\cos(\omega_{d}t)]$ in quadrature. See Appendix B for more details on the effective rotating field. In order to be deep within the adiabatic regime, we use $B_{0}=2\pi\times 80$ MHz, and modulation frequencies of $\omega_{\mathrm{mod}}=2\pi\times 0.75$ MHz.
@@ V Dissipative Stabilization of a Topological Photon Pump
Quantum states can serve as resources for classically difficult tasks, making reliable and stable preparation of these states a useful module in a variety of quantum engineering protocols. The most familiar examples of such resource states are static—a given scheme would call for the preparation of a state |$\ket{\psi_{\mathrm{resource}}}$, which does not evolve except when acted on by the protocol Giovannetti et al. (2011); Tóth and Apellaniz (2014); Briegel et al. (2009); Knill (2004); Bravyi and Kitaev (2005). However, some schemes explicitly call for time-varying periodic states, |$\ket{\psi_{\mathrm{resource}}(t)}$, and utilize this time dependence as part of their dynamical protocol Martin et al. (2017); Crowley et al. (2019); Nathan et al. (2019); Psaroudaki and Refael (2021); Long et al. (2022). These periodic resource states can be stabilized by our methods.
Cavity state boosting Long et al. (2022) is a topological pumping protocol which can be used to coherently translate, or boost, the quantum state of a cavity in the Fock basis. This enables the preparation of highly excited non-classical states, such as Fock states, if lower photon number states can be reliably prepared (see Refs. Varcoe et al. (2000); Hofheinz et al. (2008); Wang et al. (2008); Heeres et al. (2017); Rivera et al. (2023) for other methods to prepare Fock states). More generally, given an initial cavity state written in terms of Fock states
@@
The required resource state is a strongly varying qubit state that covers the Bloch sphere Martin et al. (2017); Crowley et al. (2019). When this state is chosen correctly and coupled to the cavity with exchange (Jaynes-Cummings) interactions, the influence of the qubit on the cavity is to effectively implement a Thouless pump in photon number space. On average, every cycle of the qubit state pumps a integer number $C$ of photons into the cavity, leading to an average increase in the photon number with the quantized rate
@@
where $\dot{n}=\mathrm{d}n/\mathrm{d}t$ is the average photon number current into the cavity Nathan et al. (2019). Henceforth, we take $C=1$. Further, Ref. Long et al. (2022) showed that at times $t^{*}$ which are close to being a common period of the cavity and the qubit, $t^{*}\approx N(2\pi/\Delta_{b})\approx MT_{\mathrm{mod}}$, the cavity state rephases into a boosted copy of the initial state [Eq. (12)]. The cavity frequency, $\Delta_{b}$, is defined in Eq. (14).
Reference Long et al. (2022) appealed to an adiabatic control scheme for the resource qubit. By implementing a strong, slow drive on the qubit, the back-action of the cavity on the qubit is negligible except when they need to exchange energy. This necessitates an operating regime where the pumping rate—given by $1/T_{\mathrm{mod}}$—is very low, which in turn limits the size of a cavity state that can be prepared in a time- or loss-limited experiment. The introduction of an auxiliary lossy element can stabilize the required resource state, lending additional robustness to the boosting protocol and hence allowing for faster, more reliable preparation of excited cavity states.
@@
Cavity state boosting (and topological photon pumping more broadly Nathan et al. (2019)) requires the boost cavity energy and drive frequencies to be much smaller than the qubit energy and interaction: $\Delta_{\mathrm{b}},\omega_{\mathrm{mod}}\ll B_{0},g_{\mathrm{b}}\sqrt{n_{\mathrm{b}}}$. Meanwhile, stabilization requires a resonance condition between the qubit quasienergy (which can be replaced by the instantaneous energy near the adiabatic limit) and the stabilizing cavity: $|\Delta_{\mathrm{s}}-B_{0}|\lesssim g_{\mathrm{s}}$. In order for the boost cavity to not couple strongly to the stabilizing cavity, and inherit its photon loss, we also demand that they are far detuned: $|\Delta_{\mathrm{s}}-\Delta_{\mathrm{b}}|\gg g_{\mathrm{b}},g_{\mathrm{s}}$.
@@ VI Conclusion and Outlook
However, there are two aspects of our results that suggest that the same scheme can combat these other kinds of heating. First, our experimental implementation is carried out on a transmon qubit, which is not a true two-level system. In fact, the transmon is notorious for being only loosely describable as a two-level system due to its relatively low anharmonicity of $100-250$ MHz (or roughly $2\%-5\%$). The data in Fig. 2 shows persistent oscillations for $100\ \mu$s and over $100$ drive periods without signs of uncompensated heating due to the higher levels of the transmon, despite applied drive strengths which are roughly one third of the anharmonicity. Second, the stabilized topological photon pump described in Sec. V uses a second lossy cavity to stabilize a Floquet system consisting of both a qubit and a cavity. The presence of autonomous stabilization in both of these cases despite the larger Hilbert space is a strong indicator that it is possible to suppress Floquet heating in general using our simple scheme, rather than the contributions of excited-state lifetime alone. There is also evidence that the scheme can combat heating due to many-body interactions, and can cool into desired ground states of effective Floquet Hamiltonians Petiziol and Eckardt (2022); Schaefer et al. . However, a full characterization of the capability of autonomous stabilization with a static auxiliary remains an outstanding question.
Optimal stabilization, with the fastest possible stabilization time and highest possible steady-state purity, relies on exact resonances in the augmented spin-cavity system. However, significant stabilization can be achieved for a much wider range of parameters. Supplementary experimental data, shown in Appendix F, shows that significant stabilization is observed for all configurations in the range $|\delta|\lesssim g$. As the transmon decay rate is only a fourth of the photon decay rate, using an even lossier cavity (up to $\kappa\approx g$) would result in stabilization for an even wider range of parameters. Additionally, we find that elliptical drive fields, which are only resonant during part of the drive cycle, also exhibit stabilization. The combination of these two experimental observations is a strong indicator that a single lossy auxiliary degree of freedom can be used to counteract heating into all excited states within a range of roughly $g$ of the resonance, and hence systematically remove population from entire targeted Floquet bands. Alternatively, an auxiliary consisting of a band-pass filter, rather than a single cavity, could be used to enhance loss for a large number of Floquet states Wang and Safavi-Naeini (2023). Implementations along these lines would give the ability to, for example, cool directly into the ground band of a Floquet engineered lattice Seetharam et al. (2015), rather than relying on adiabatic band mapping, which is generally only possible in synthetic lattice systems.